# Robust Equilibria in Shared Resource Allocation via Strengthening Border’s Theorem

### v1

@@ Appendix C Border’s Theorem
$\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}]=\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right)=\beta_{i}p_{i}.$ (15)
We observe that $p_{i}^{S}=0$ if $i\notin S$, and so by definition, the $p_{i}^{S}$’s induce the interim allocation probabilities $p_{i}$. ∎
C.3 A Flow Network Proof of Theorem C.2
A proof of Border’s Theorem based on flow networks and the max-flow min-cut theorem was discovered by [che2013generalized]. We will give flow network proof in the special case of Border’s Theorem that we use (Theorem C.2) here for convenience.
@@
Figure 4: Flow network that can be used to prove Theorem A.3. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\beta_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem A.3 by analyzing every minimum-cut of this network.
Flow network proof of Theorem C.2.
As argued in the remarks after Theorem C.2, (10) is necessary for the $p_{i}^{S}$’s to exist, so we assume it.
@@
$c(v_{i},t)=p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime}).$ (17)
The flow network is depicted in Fig. 4.
@@
$\sum_{S\subseteq[n]:S\neq\emptyset}c(s,u_{S})=\sum_{S\subseteq[n]:S\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\neq\emptyset)=1-\prod_{i\in S}(1-\beta_{i}).$ (18)
@@
$\sum_{i\in[n]}c(v_{i},t)=\sum_{i\in[n]}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})=\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i}),$ (19)
using (10) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (11) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
$\begin{split}c(A,B)&=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}c(s,u_{S})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)+\sum_{i\notin I}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})\\ &=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\cap I\neq\emptyset)+\sum_{i\notin I}p_{i}\beta_{i}\\ &=1-\prod_{i\in I}(1-\beta_{i})+1-\prod_{i\in[n]}(1-\beta_{i})-\sum_{i\in I}p_{i}\beta_{i},\end{split}$ (20)
using (10) for the last line. Rearranging, one can see that (11) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$. ∎
C.4 Theorem C.2 with Arbitrary Upper Bounds
@@
$\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$ (21)
and for any $I\subseteq[n]$ and $\mathcal{S}\subseteq\{S\subseteq[n]:S\cap I\neq\emptyset\}$,
$\sum_{i\in I}p_{i}\beta_{i}+\prod_{i\in I}(1-\beta_{i})+\sum_{S\in\mathcal{S}}\left(\prod_{i\in S}\beta_{i}\right)\left(\prod_{i\notin S}(1-\beta_{i})\right)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)\leq 1.$ (22)
Proof.
Since (21) is necessary for the $p_{i}^{S}$’s to exist in Theorem 5.2, it is also necessary here, so we assume it.
@@
The flow network is depicted in Fig. 5.
@@
Figure 5: Flow network that is used in the proof of Theorem C.4. The flow network is the same as Fig. 4 except the middle edges $(u_{S},v_{i})$ have capacity $\bar{p}_{i}^{S}\Pr(S^{\prime}=S)$ to enforce the upper bounds on the allocation probabilities.
@@
using (21) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{S}\leq\bar{p}_{i}^{S}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (22) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
using (21) for the last line. Rearranging, (22) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$. ∎
Appendix D Deferred Proofs
D.1 Deferred Proofs of Corollaries A.4 and A.5
The proof of Corollary A.4 is almost identical to proof of Theorem 4.2 given in Section 5.2.
Proof of Corollary A.4.
We prove that the choice of $p_{i}=\frac{1-\prod_{j=1}^{n}(1-\gamma\alpha_{j})}{\gamma}$ satisfies the conditions of Theorem A.3. Observe that (8) holds with this choice of $p_{i}$.
Also, observe that (3) holds for $I=\emptyset$. To show that (3) holds for $I\neq\emptyset$, define $a_{I}$ as
@@ Appendix D Deferred Proofs
With this choice of $p_{i}$, (3) says
@@
This holds if and only if $a_{I}\geq\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. It suffices to show $a_{I}$ is nonincreasing in $I$ since $a_{[n]}=\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. To do this, we can simply follow the same proof that $a_{I}$ as defined in the proof of Theorem 4.2 is nonincreasing, but replace each $\alpha_{i}$ with $\beta_{i}$. We did not use the fact that $\sum_{i=1}^{n}\alpha_{i}=1$ for that part of the proof of Theorem 4.2, so everything still works. ∎
The proof of Corollary A.5 is an application of Corollary A.4.
Proof of Corollary A.5.
@@
Using Corollary A.4, we can find interim allocation probabilities $p_{i}$ such that,
@@
The result follows from substituting $\alpha=\frac{1}{n}$ and $\gamma=\Theta(\log n)$. ∎
D.2 Deferred Proofs of Lemma 5.1, Propositions 4.1 and A.1, Theorems 4.3 and A.2
We shall prove these in the context of our more general mechanism Generalized Budgeted Border, which is the same as Budgeted Robust Border except the each agent gets $\beta_{i}(1+\delta_{i}^{T})$ budget of bid tokens where $\beta_{i}\in[0,1]$ is a parameter. Budgeted Robust Border is Generalized Budgeted Border when each $\beta_{i}=\alpha_{i}$. See Appendix A for details.
@@
Lemma D.1.
Fix an agent $i$ and assume all other agents $j\neq i$ are playing in a way such that the sets $S_{\neq i}[t]=\{j\neq i:b_{j}^{t}=1\}$ of bidding agents $j\neq i$ are i.i.d. across rounds drawn from some distribution $\nu$ over subsets of $[n]\setminus\{i\}$. Suppose agent $i$ is trying to maximize her imaginary expected utility subject to the constraint that she does not exceed her budget in expectation; that is, she is choosing $(\tilde{b}_{i}^{t})$ to solve
$\max\,\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]\quad\text{subject to}\quad\sum_{t=1}^{T}\mathbb{E}[\tilde{b}_{i}^{t}]\leq\beta_{i}^{\prime}T$ (23)
@@
Notice that Lemma 5.1 is a special case of the above lemma with parameters $\beta_{i}=\alpha_{i}$, $\delta_{i}^{T}=0$, and $\nu$ being the distribution over subsets $S_{\neq i}$ where each $j\in S_{\neq i}$ independently with probability $\alpha_{j}$. In the proof, we use the notion of $\beta$-ideal utility and $\beta$-ideal utility probability function from Definition A.1 defined in Appendix A.
Proof of Lemma D.1.
@@
$\begin{split}\mathbb{E}[\tilde{U}_{i}[t]]=\mathbb{E}[V_{i}[t]\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}]]\\ =\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]\mathbb{E}[\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}=1]=p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}].\end{split}$ (25)
Thus, agent $i$’s maximization problem is equivalent to maximizing $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$ subject to the same constraint in (23). It is clear for any feasible solution $(\tilde{b}_{i}^{t})$ that the solution that bids at time $t$ with probability $\rho[t](V_{i}[t])=\Pr(\tilde{b}_{i}^{t}=1\mid V_{i}[t])$ is also a feasible solution with the same objective value. Thus, we can rewrite the agent’s maximization problem in terms of maximizing over $\rho[t](V_{i}[t])$, i.e., the following maximization problem over measurable functions $\rho[t]:[0,\infty)\to[0,1]$:
@@
Given any optimal solution $(\rho[t])$ to the above, observe that setting $\rho^{\star}[t]=\frac{1}{T}\sum_{s=1}^{T}\rho[s]$ is also a feasible solution with the same objective value. Observe then that $\mathbb{E}[\rho^{\star}[t](V_{i}[t])]=\beta_{i}^{\prime}$, and so $\rho^{\star}[t]$ is a feasible solution in (A.1), the definition of $\beta_{i}^{\prime}$-ideal utility, and it must maximize the same objective $\mathbb{E}[V_{i}[t]\rho[t](V_{i}[t])]]$. Therefore, $\rho^{\star}[t]$ is exactly the $\beta_{i}^{\prime}$-ideal utility probability function $(\rho_{i}^{\beta_{i}^{\prime}})^{\star}$, so the optimal bidding strategy $(\tilde{b}_{i}^{t})^{\star}$ to solve (23) is precisely a $\beta_{i}^{\prime}$-aggressive strategy.
By (25), under such a $\beta_{i}^{\prime}$-aggressive strategy, agent $i$ obtains utility $p_{i}(\nu)v^{\star}(\beta_{i}^{\prime})$. ∎
@@
The result follows from substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$. ∎
@@
Lemma D.3.
Fix an agent $i$ playing a $\beta_{i}$-aggressive strategy. Suppose the other agents $j\neq i$ are playing in the imaginary game with strategies as in Lemma D.1. Let $E_{1}$ be an event on which agents $j\neq i$ do not run out of budget in the actual game. Then, on an subevent of $E_{1}$ with probability at least $\Pr(E_{1})-O(1/T^{2})$,
@@
Proof.
Observe that the random variables $\tilde{U}_{i}[t]=V_{i}[t]\boldsymbol{1}\{\tilde{i}^{t}=i\}$ are i.i.d. By (25), each has mean $p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$. Since player $i$ is playing a $\beta_{i}$-aggressive strategy, this mean is $\mathbb{E}[\tilde{U}_{i}[t]]=p_{i}(\nu)v_{i}^{\star}(\beta_{i})$. Recall that we assume the value distribution $\mathcal{F}_{i}$ is bounded so the $U_{i}$ are bounded by some $\bar{v}$. Let $\epsilon>0$. By Hoeffding’s inequality,
$\Pr\left(\left|\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})T\right|\geq\epsilon\right)\leq 2\exp\left(-\frac{2\epsilon^{2}}{\bar{v}^{2}T}\right)$ (26)

### v2

@@ D Border’s Theorem
(D.6) $\operatorname*{\mathbb{E}}_{S\sim\mu}[p_{i}^{S}]=\alpha_{i}\left(1-\prod_{j=1}^{n}(1-\beta_{j})\right)=\beta_{i}p_{i}.$
We observe that $p_{i}^{S}=0$ if $i\notin S$, and so by definition, the $p_{i}^{S}$’s induce the interim allocation probabilities $p_{i}$.
That was the proof for the case where (D.2) holds with strict inequality; we now give a topological argument to extend to the general case. Let $B$ be the set of bid rates and interim allocation probabilities $(\beta_{i},p_{i})$ that satisfy Border’s Criterion. In subsection˜D.2, we showed that the set of $(\beta_{i},p_{i})\in B$ satisfying (B.2) with strict inequality for $I\subsetneq[n]$ are induced by allocation probabilities.
is compact being the inverse image of a closed set under a continuous mapping, and the set of $(\beta_{i},p_{i})$ that are induced by allocation probabilities is just the projection of this set onto the first two coordinates. Hence, it suffices to show that the set of $(\beta_{i},p_{i})\in B$ satisfying (B.2) with strict inequality for $I\subseteq[n]$ is dense in $B$. Given any such $(\beta_{i},p_{i})\in B$, letting $\varepsilon>0$, define $(\beta_{i}^{\prime},p_{i}^{\prime})$ by
$\beta_{i}^{\prime}$ $=\beta_{i}+\varepsilon$
Then, $(\beta_{i}^{\prime},p_{i}^{\prime})$ can be made arbitrarily close to $(\beta_{i},p_{i})$. They can be seen to satisfy Border’s Criterion where (B.2) is strict for $I\subsetneq[n]$ as follows. We have increased $\beta_{i}$ and decreased $p_{i}$ to get $(\beta_{i}^{\prime},p_{i}^{\prime})$ in such a way that both the left-hand side and right-hand side of (B.1) increase in the same amount to maintain equality. In (B.2) for $I\subsetneq[n]$, considering the changes caused by each coordinate $i\in I$ one at a time, the left-hand side increases by the same amount as in (B.1) but the right-hand side increases by a larger amount since the partial derivative of the right-hand side with respect to some $\beta_{i}\in I$ is (strictly) decreasing in $I$ (when all $\beta_{j}>0$ for $j\in I$). Therefore, with the $(\beta_{i}^{\prime},p_{i}^{\prime})$, (B.2) holds with strict inequality.
D.3 A Flow Network Proof of Theorem˜D.2
A proof of Border’s Theorem based on flow networks and the max-flow min-cut theorem was discovered by [che2013generalized]. We will give flow network proof in the special case of Border’s Theorem that we use (Theorem˜D.2) here for convenience.
@@
Figure D.1: Flow network that can be used to prove Theorem˜B.5. We let $S^{\prime}$ be the random set of bidding agents where each agent $i$ lies in $S^{\prime}$ independently with probability $\beta_{i}$. Then, there are three kind of edges: edges whose flow corresponds to the probability of observing a specific $S^{\prime}$ (left), edges whose flow corresponds to how we randomly allocate the item condition on observing a specific $S^{\prime}$ (middle), and edges whose flow represent the probability that a specific agent gets the item (right). There is a flow of value $\Pr(S^{\prime}\neq\emptyset)$ if and only if there exist allocation probabilities $p_{i}^{S}$ inducing the interim allocation probabilities $p_{i}$. In other words, the flows $p_{i}^{S}\Pr(S^{\prime}=S)$ in the middle transform the probabilities that agents in a certain set $S$ bid to an agent $i\in S$ being allocated. We obtain the conditions in Theorem˜B.5 by analyzing every minimum-cut of this network.
Proof 7 (Flow network proof of Theorem˜D.2).
As argued in the remarks after Theorem˜D.2, (D.1) is necessary for the $p_{i}^{S}$’s to exist, so we assume it.
@@
(D.8) $c(v_{i},t)=p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime}).$
The flow network is depicted in Figure˜D.1.
@@
(D.9) $\sum_{S\subseteq[n]:S\neq\emptyset}c(s,u_{S})=\sum_{S\subseteq[n]:S\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\neq\emptyset)=1-\prod_{i\in S}(1-\beta_{i}).$
@@
(D.10) $\sum_{i\in[n]}c(v_{i},t)=\sum_{i\in[n]}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})=\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i}),$
using (D.1) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (D.2) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
(D.11) $\begin{split}c(A,B)&=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}c(s,u_{S})+\sum_{i\notin I}c(v_{i},t)\\ &=\sum_{S\subseteq[n]:S\cap I\neq\emptyset}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)+\sum_{i\notin I}p_{i}\Pr_{S^{\prime}\sim\mu}(i\in S^{\prime})\\ &=\Pr_{S^{\prime}\sim\mu}(S^{\prime}\cap I\neq\emptyset)+\sum_{i\notin I}p_{i}\beta_{i}\\ &=1-\prod_{i\in I}(1-\beta_{i})+1-\prod_{i\in[n]}(1-\beta_{i})-\sum_{i\in I}p_{i}\beta_{i},\end{split}$
using (D.1) for the last line. Rearranging, one can see that (D.2) is equivalent to the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$.
D.4 Theorem˜D.2 with Arbitrary Upper Bounds
@@
(D.12) $\sum_{i\in[n]}p_{i}\beta_{i}=1-\prod_{i\in[n]}(1-\beta_{i})$
and for any $I\subseteq[n]$,
$\sum_{i\in I}p_{i}\beta_{i}+\prod_{i\in I}(1-\beta_{i})+\sum_{S\in\mathcal{S}}\left(\prod_{i\in S}\beta_{i}\right)\left(\prod_{i\notin S}(1-\beta_{i})\right)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)\leq 1$
where $\mathcal{S}=\{S\subseteq[n]:S\cap I\neq\emptyset,\sum_{i\in S\cap I}\bar{p}_{i}^{S}\leq 1\}$.
Proof 8.
Since (D.12) is necessary for the $p_{i}^{S}$’s to exist in Theorem˜4.2, it is also necessary here, so we assume it.
@@
The flow network is depicted in Figure˜D.2.
@@
Figure D.2: Flow network that is used in the proof of Theorem˜D.4. The flow network is the same as Figure˜D.1 except the middle edges $(u_{S},v_{i})$ have capacity $\bar{p}_{i}^{S}\Pr(S^{\prime}=S)$ to enforce the upper bounds on the allocation probabilities.
@@
using (D.12) for the last equality. Observe that in this flow network, allocation probabilities $(p_{i}^{S})$ satisfying upper bounds $p_{i}^{S}\leq\bar{p}_{i}^{S}$ induce the interim allocation probabilities $(p_{i})$ if and only if the flow $f$ is feasible where $f(s,u_{S})=c(s,u_{S})$, $f(v_{i},t)=c(v_{i},t)$, and $f(u_{S},v_{i})=p_{i}^{S}\Pr_{S^{\prime}\sim\mu}(S^{\prime}=S)$. Since both the $s$-$t$ cuts with $s$ on one side and everything else on the other and the cut with $t$ on one side and everything else on the other both have cut capacity $1-\prod_{i\in[n]}(1-\beta_{i})$, it suffices to show that (D.4) holds only if there is a feasible flow of flow value equal to this cut capacity.
@@
using (D.12) for the last line. Rearranging, the above being least $1-\prod_{i\in[n]}(1-\beta_{i})$ is equivalent to
$\sum_{i\in I}p_{i}\beta_{i}+\prod_{i\in I}(1-\beta_{i})+\sum_{S\in\mathcal{S}}\left(\prod_{i\in S}\beta_{i}\right)\left(\prod_{i\notin S}(1-\beta_{i})\right)\left(1-\sum_{i\in S\cap I}\bar{p}_{i}^{S}\right)\leq 1.$
Thus, the flow network is feasible if and only if the above holds for all $I\subseteq[n]$ and $\mathcal{S}\subseteq\{S\subseteq[n]:S\cap I\neq\emptyset\}$. To maximize the left-hand side above, we should set $\mathcal{S}=\{S\subseteq[n]:S\cap I\neq\emptyset,\sum_{i\in S\cap I}\bar{p}_{i}^{S}\leq 1\}$, which gives the theorem statement.
E Deferred Proofs
E.1 Deferred Proof of Lemma˜2.2
Proof 9 (Proof of Lemma˜2.2).
(E.1) $\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\leq 1-\prod_{j\in[n]}(1-\alpha_{j}).$
For these value distributions, agent $i$’s ideal utility is $\alpha_{i}$. If each agent obtains a $\lambda$ fraction of their ideal utility in expectation, then $\frac{1}{T}\sum_{t=1}^{T}\sum_{i\in[n]}\mathbb{E}[U_{i}[t]]\geq\lambda\sum_{i\in[n]}\alpha_{i}=\lambda,$ so $\lambda\leq 1-\prod_{j\in[n]}(1-\alpha_{j})$ by (E.1).
E.2 Deferred Proofs of Theorems˜3.5, B.6, and B.7
Proof 10 (Proof of Theorem˜3.5).
We prove that the choice of $p_{i}=1-\prod_{j=1}^{n}(1-\alpha_{j})$ satisfies the conditions of Theorem˜4.2. Observe that (4.1) holds with equality when $I=[n]$ for this choice of $p_{i}$. Also, observe that (4.1) holds for $I=\emptyset$. To show that (4.1) holds for $I\neq\emptyset$, define $a_{I}$ as
$a_{I}=\frac{1-\prod_{i\in I}(1-\alpha_{i})}{\sum_{i\in I}\alpha_{i}}.$
With this choice of $p_{i}$, (4.1) says
This can be seen to be nonpositive by the identities that $1-x\leq e^{-x}$, so $\prod_{i\in I\setminus\{k\}}(1-\alpha_{i})\leq\exp\left(-\sum_{i\in I\setminus\{k\}}\alpha_{i}\right)$, and also $e^{-y}(1+y)\leq 1$ applied to $y=\sum_{i\in I\setminus\{k\}}\alpha_{i}$.
The proof of Corollary˜B.6 is almost identical to proof of Theorem˜3.5.
Proof 11 (Proof of Corollary˜B.6).
We prove that the choice of $p_{i}=\frac{1-\prod_{j=1}^{n}(1-\gamma\alpha_{j})}{\gamma}$ satisfies the conditions of Theorem˜B.5. Observe that (B.1) holds with this choice of $p_{i}$.
Also, observe that (B.2) holds for $I=\emptyset$. To show that (B.2) holds for $I\neq\emptyset$, define $a_{I}$ as
@@ E Deferred Proofs
With this choice of $p_{i}$, (B.2) says
@@
This holds if and only if $a_{I}\geq\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. It suffices to show $a_{I}$ is nonincreasing in $I$ since $a_{[n]}=\frac{1-\prod_{j=1}^{n}(1-\beta_{j})}{\gamma}$. To do this, we can simply follow the same proof that $a_{I}$ as defined in the proof of Theorem˜3.5 is nonincreasing, but replace each $\alpha_{i}$ with $\beta_{i}$. We did not use the fact that $\sum_{i=1}^{n}\alpha_{i}=1$ for that part of the proof of Theorem˜3.5, so everything still works.
The proof of Corollary˜B.7 is an application of Corollary˜B.6.
Proof 12 (Proof of Corollary˜B.7).
@@
Using Corollary˜B.6, we can find interim allocation probabilities $p_{i}$ such that,
@@
The result follows from substituting $\alpha=\frac{1}{n}$ and $\gamma=\Theta(\log n)$.
E.3 Deferred Proofs of Lemma˜4.1, Propositions˜3.3 and B.2, Theorems˜3.6 and B.4
We shall prove these in the context of our more general mechanism Generalized Budgeted Border, which is the same as Budgeted Robust Border except the each agent gets $\beta_{i}(1+\delta_{i}^{T})$ budget of bid tokens where $\beta_{i}\in[0,1]$ is a parameter. Budgeted Robust Border is Generalized Budgeted Border when each $\beta_{i}=\alpha_{i}$. See section˜B for details.
@@
Lemma E.1.
Fix an agent $i$ and assume all other agents $j\neq i$ are playing in a way such that the sets $S_{\neq i}[t]=\{j\neq i:b_{j}^{t}=1\}$ of bidding agents $j\neq i$ are i.i.d. across rounds drawn from some distribution $\nu$ over subsets of $n]\setminus\{i\}$. Suppose agent $i$ is trying to maximize her imaginary expected utility subject to the constraint that she does not exceed her budget in expectation; that is, she is choosing $(\tilde{b}_{i}^{t})$ to solve
(E.2) $\max\,\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\tilde{U}_{i}[t]]\quad\text{subject to}\quad\sum_{t=1}^{T}\mathbb{E}[\tilde{b}_{i}^{t}]\leq\beta_{i}^{\prime}T$
@@
Notice that Lemma˜4.1 is a special case of the above lemma with parameters $\beta_{i}=\alpha_{i}$, $\delta_{i}^{T}=0$, and $\nu$ being the distribution over subsets $S_{\neq i}$ where each $j\in S_{\neq i}$ independently with probability $\alpha_{j}$. In the proof, we use the notion of $\beta$-ideal utility and $\beta$-ideal utility probability function from Definition˜B.1 defined in section˜B.
Proof 13 (Proof of Lemma˜E.1).
@@
(E.4) $\begin{split}\mathbb{E}[\tilde{U}_{i}[t]]=\mathbb{E}[V_{i}[t]\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}]=\mathbb{E}[\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}]]\\ =\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]\mathbb{E}[\boldsymbol{1}\{i=\tilde{i}^{t}\}\mid\tilde{b}_{i}^{t}=1]=p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}].\end{split}$
Thus, agent $i$’s maximization problem is equivalent to maximizing $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$ subject to the same constraint in (E.2). It is clear for any feasible solution $(\tilde{b}_{i}^{t})$ that the solution that bids at time $t$ with probability $\rho[t](V_{i}[t])=\Pr(\tilde{b}_{i}^{t}=1\mid V_{i}[t])$ is also a feasible solution with the same objective value. Thus, we can rewrite the agent’s maximization problem in terms of maximizing over $\rho[t](V_{i}[t])$, i.e., the following maximization problem over measurable functions $\rho[t]:[0,\infty)\to[0,1]$:
@@
Given any optimal solution $(\rho[t])$ to the above, observe that setting $\rho^{\star}[t]=\frac{1}{T}\sum_{s=1}^{T}\rho[s]$ is also a feasible solution with the same objective value. Observe then that $\mathbb{E}[\rho^{\star}[t](V_{i}[t])]=\beta_{i}^{\prime}$, and so $\rho^{\star}[t]$ is a feasible solution in (B.1), the definition of $\beta_{i}^{\prime}$-ideal utility, and it must maximize the same objective $\mathbb{E}[V_{i}[t]\rho[t](V_{i}[t])]]$. Therefore, $\rho^{\star}[t]$ is exactly the $\beta_{i}^{\prime}$-ideal utility probability function $(\rho_{i}^{\beta_{i}^{\prime}})^{\star}$, so the optimal bidding strategy $(\tilde{b}_{i}^{t})^{\star}$ to solve (E.2) is precisely a $\beta_{i}^{\prime}$-aggressive strategy.
By (E.4), under such a $\beta_{i}^{\prime}$-aggressive strategy, agent $i$ obtains utility $p_{i}(\nu)v^{\star}(\beta_{i}^{\prime})$.
@@
The result follows from substituting $\delta_{i}^{T}=\sqrt{\frac{6\ln T}{\beta_{i}T}}$.
@@
Lemma E.3.
Fix an agent $i$ playing a $\beta_{i}$-aggressive strategy. Suppose the other agents $j\neq i$ are playing in the imaginary game with strategies as in Lemma˜E.1. Let $E_{1}$ be an event on which agents $j\neq i$ do not run out of budget in the actual game. Then, on an subevent of $E_{1}$ with probability at least $\Pr(E_{1})-O(1/T^{2})$,
@@
Proof 15.
Observe that the random variables $\tilde{U}_{i}[t]=V_{i}[t]\boldsymbol{1}\{\tilde{i}^{t}=i\}$ are i.i.d. By (E.4), each has mean $p_{i}(\nu)\mathbb{E}[V_{i}[t]\tilde{b}_{i}^{t}]$. Since player $i$ is playing a $\beta_{i}$-aggressive strategy, this mean is $\mathbb{E}[\tilde{U}_{i}[t]]=p_{i}(\nu)v_{i}^{\star}(\beta_{i})$. Recall that we assume the value distribution $\mathcal{F}_{i}$ is bounded so the $U_{i}$ are bounded by some $\bar{v}$. Let $\varepsilon>0$. By Hoeffding’s inequality,
(E.5) $\Pr\left(\left|\sum_{t=1}^{T}\tilde{U}_{i}[t]-p_{i}(\nu)v_{i}^{\star}(\beta_{i})T\right|\geq\varepsilon\right)\leq 2\exp\left(-\frac{2\varepsilon^{2}}{\bar{v}^{2}T}\right)$